# The accreted stellar haloes of Milky Way-mass galaxies as a probe of the nature of the dark matter

### v1

@@ 4.4 The building blocks
To find which galaxies contributed to the build up of the stellar haloes in our sample, we identify all luminous structures whose main merger tree branches directly merge with the main progenitor of each MW-mass halo. We show in Fig. 5 the resulting $V_{\mathrm{peak}}$ [^4] Defined as the peak value of the maximum circular velocity of a subhalo, $V_{\mathrm{circ}}=\sqrt{GM(<r)/r}$, as measured across its main progenitor branch. function of the population we find, which is a proxy for halo mass, and hence stellar mass.
@@
The differences discussed in §4.2 and §4.3 concern the inner region of the stellar halo, and so are likely to be driven by its most massive progenitors. However, no significant differences are present in this mass range across models, which are all consistent within their 16th - 84th percentiles. As the models we have used here primarily affect their internal structure, or the masses and abundances of less massive haloes, the total masses of the largest contributors are unaffected, and how much mass they deposit remains similar.
@@
To do this, we integrate the orbits in the particle potential around the MW-mass main progenitor. The potential is fit to the particle distribution using AGAMA (Vasiliev, 2019) and assuming a spherically symmetric potential for the stars, gas and dark matter particles. We believe that this assumption, though approximate, is sufficient to capture the major trends in the population we study here. As initial conditions, we use the relative position and velocity between the stellar halo progenitors and the MW-mass main progenitor, measured at the output when the former were last identified as self-bound by SUBFIND.
The resulting distribution of pericentric distances estimated at the time before disruption is shown in Fig. 6. Since the mass of the progenitor is an important indicator for whether they dominate the mass budget of the stellar halo, we split the distribution into two based on whether their $V_{\mathrm{peak}}$ is greater or lower than $50~{}\mathrm{km}\,\mathrm{s}^{-1}$. We also show these distributions normalised to the total number of stellar halo progenitors, to decouple the suppression of structure in WDM from the distribution of their disruption distances.
As the overall properties of the stellar halo are set by those with the highest $V_{\mathrm{peak}}$ values, we focus first on the $V_{\mathrm{peak}}>50~{}\mathrm{km}\,\mathrm{s}^{-1}$ bin. Despite a similar number of progenitors across DM models, the shape of the distributions are very different. The clearest change occurs in SIDM, where most objects disrupt at larger pericentric distances than CDM and WDM. Since the initial orbital properties are expected to be similar to CDM, the differences arise due to the subsequent orbital evolution and mass loss that the satellite galaxies experienced. As argued previously, the faster mass loss rates lead to an earlier disruption, and hence potentially lessened effects of dynamical friction and radialisation. This would generally result in their stellar remnants retaining higher orbital energy and being less radially supported than in CDM, as discussed in §4.2 and §4.3.
@@ 5 Conclusions
  • The spatial distribution of stellar haloes is more sensitive to the assumed model of dark matter, and are largely present in the inner parts (Fig. 2). This is most evident for the SIDM model, where the differences already appear within $0.05R_{200}$ ($\sim 10\,\mathrm{kpc}$), coinciding with the spatial scale where the DM density profile has flattened due to self-interactions. The changes manifest as a reduction of the accreted stellar density to a median value of $\langle\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}\rangle=0.3^{+0.9}_{-0.2}$ of its CDM counterparts, as well as shallower density profiles: $\langle\alpha_{\mathrm{SIDM}}\rangle=-1.4^{+0.4}_{-0.3}$ vs $\langle\alpha_{\mathrm{CDM}}\rangle=-1.7^{+0.3}_{-0.4}$. In WDM, the profiles change in the opposite direction than SIDM, being instead slightly denser, $\langle\rho^{\mathrm{WDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}\rangle=1.4^{+0.6}_{-0.6}$, and steeper, $\langle\alpha_{\mathrm{WDM}}\rangle=-2.0^{+0.1}_{-0.4}$ than the CDM counterparts. However, the changes in the WDM ex-situ density are minor compared to those observed in the SIDM model.
  • Beyond their density profiles, the velocity distribution of accreted stars also differs across models. This results in a larger number and prominence of overdensities in $r,v_{r}$ space (Fig. 3) and more tangentially supported orbits in SIDM than in CDM (Fig. 4). This amounts to a typical change in the velocity anisotropy of $\langle\Delta\beta\rangle=-0.2^{+0.1}_{-0.1}$ between SIDM and CDM in the inner $0.01R_{200}$ ($\sim 2\,\mathrm{kpc}$). The changing anisotropy could be used in conjunction with the properties of the GE/S in the Milky Way to place upper limits on the allowed SIDM cross-section. In other words, what is the maximum possible value of $\sigma_{\mathrm{SIDM}}$ that can accommodate highly radially supported ($\beta\sim 0.8$) remnant of GE/S-like events. This will be explored in upcoming work.
@@
Several assumptions and limitations are present in this study. Firstly, we have considered simulations in which the predicted differences between dark matter models are exacerbated, i.e. no flat-density core formation through gas blowouts in dwarf galaxies. As shown in Forouhar Moreno et al. (2022), this process can mimic the effects of SIDM on the satellite population, and could therefore similarly affect the most massive stellar halo progenitors and thus its final properties. It nonetheless remains to be seen whether baryon-driven outflows are effective at forming flat density cores for massive GE/S-like galaxies at high redshifts. Secondly, we have used relatively extreme variations in the nature of dark matter, and so follow-up work should consider versions that are more consistent with current observational constraints. Lastly, our comparisons between dark matter models do not include observational uncertainties. This means that the differences we identified in this work may be difficult to observe in the real Universe. As such, mock observations should be made based on these simulations, to enable a fairer comparison to real data.
@@
Despite these caveats, our findings highlight a new potential avenue to constrain the nature of dark matter, based on the dynamics and distribution of the accreted stellar halo of Milky Way-mass galaxies. One approach would be to measure the spatial distribution of the stellar halo around a large sample of MW-like galaxies in the Universe, and compare it to the predicted distribution across different DM models. An alternative would be to leverage the phase-space information of GE/S, by identifying simulated MW-analogues that experienced similar events in their past and measuring the final velocity anisotropy of its remnant. As changing the value of the SIDM cross-section systematically affects the merging process, and hence the anisotropy $\beta$, comparing the predicted distributions of $\beta$ to the one observed for GE/S could hint at which SIDM cross-sections are compatible with the data.
@@
Beyond these two approaches, one can imagine other observables which we have not considered and could be similarly affected, such as the two-point correlation function in phase-space, clustering in action space, and metallicity gradients. Given ongoing and upcoming efforts towards studying the stellar halo Milky Way and those of external Milky Way-mass haloes, this exciting prospect warrants further investigation.
@@ Acknowledgements
VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025) and the European Research Council (ERC) through the Advanced Investigator grant DMIDAS (GA 786910) and Consolidated Grant ST/T000244/1. A.B.L. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (GA 101026328). AD is supported by a Royal Society University Research Fellowship and the Leverhulme Trust. AD and AF acknowledge support from the Science and Technology Facilities Council (STFC) [grant numbers ST/X001075/1, ST/T000244/1]. AF is supported by a UKRI Future Leaders Fellowship (grant no MR/T042362/1). This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@ Appendix B Radial velocity vs radial distance of accreted stars
Figure 8: Radial velocity of accreted stellar particles, as a function of their radial distance to the centre of their host galaxies. The colour of each particle encodes the local ex-situ stellar density. This is shown for the remainder of the haloes not shown in Fig. 3, with their their matched CDM, WDM and SIDM versions placed from left to right. The vertical ordering reflects the $M_{200}(z=0)$ ranking, with decreasing mass towards the lower panels.

### v2

@@ 4.4 The properties of stellar halo progenitors
To find which galaxies contributed to the build up of the stellar haloes in our sample, we identify all luminous structures whose main merger tree branch directly merges with the main progenitor branch of each MW-mass halo. We show in Fig. 5 the resulting $V_{\mathrm{peak}}$ [^7] Defined as the peak value of the maximum circular velocity of a subhalo, $V_{\mathrm{circ}}=\sqrt{GM(<r)/r}$, as measured across its main progenitor branch. function of the population we find, which is a proxy for halo mass, and hence (approximately) stellar mass.
@@
The differences discussed in §4.2 and §4.3 concern the inner region of the stellar halo, and so are likely to be driven by its most massive progenitors. However, no significant differences are present in this mass range across models, which are all consistent within their 16th - 84th percentiles, with the most massive progenitor typically reaching $V_{\rm peak}\approx 100\,\mathrm{km}\,\mathrm{s}^{-1}$. As the models we have used here primarily affect their internal structure, or the masses and abundances of less massive haloes, the total masses of the largest contributors are unaffected, and how much mass they deposit remains similar.
@@
To do this, we integrate the orbits of all galaxies that are progenitors of the stellar halo in the particle potential around their corresponding MW-mass main progenitor. The potential is fit to the particle distribution using AGAMA (Vasiliev, 2019) and assuming a spherically symmetric potential for the stars, gas and dark matter particles. We believe that this assumption, though approximate, is sufficient to capture the major trends in the population we study here. As initial conditions, we use the relative position and velocity between the stellar halo progenitors and the MW-mass main progenitor, measured at the output when the former were last identified as self-bound by SUBFIND. We perform this exercise for each of the eight MW-mass galaxies in our sample to obtain the corresponding distribution of pericentric distances as measured before the progenitors of their stellar haloes disrupted.
The distribution of pericentric distances just before disruption, based on the median of the distributions obtained from the eight MW-mass galaxies in our sample, is shown in Fig. 6. Since the mass of the progenitor is an important indicator for whether they dominate the mass budget of the stellar halo, we split the distribution into two based on whether their $V_{\mathrm{peak}}$ is greater or lower than $50~\mathrm{km}\,\mathrm{s}^{-1}$. We also show these distributions normalised to the total number of stellar halo progenitors, to decouple the suppression of structure in WDM from the distribution of their disruption distances.
As the overall properties of the stellar halo are set by those with the highest $V_{\mathrm{peak}}$ values, we focus first on the $V_{\mathrm{peak}}>50~\mathrm{km}\,\mathrm{s}^{-1}$ bin. Despite a similar number of progenitors across DM models, the shape of the distributions are very different. The clearest change occurs in SIDM, where most objects disrupt at larger pericentric distances than CDM and WDM. Since the initial orbital properties are expected to be similar to CDM, the differences arise due to the subsequent orbital evolution and mass loss that the satellite galaxies experienced. As argued previously, the faster mass loss rates lead to an earlier disruption, and hence potentially lessened effects of dynamical friction and radialisation. This would generally result in their stellar remnants retaining higher orbital energy and being less radially supported than in CDM, as was found in §4.2 and §4.3. Despite this argument being particularly true for G1, G2 and G3, we have no reasons to believe this would not be the case in general.
@@ 5 Conclusions
  • The spatial distribution of stellar haloes is more sensitive to the assumed model of dark matter than their ex-situ stellar mass. The differences caused by different dark matter models become the most prominent towards the inner stellar halo, with the outer stellar halo showing little-to-no differences (Fig. 2). In the SIDM model, the differences appear within $0.05R_{200}$ ($\approx 10\,\mathrm{kpc}$), coinciding with the spatial scale where the DM density profile has flattened due to self-interactions. The changes manifest as a reduction of the accreted stellar density to a median value of $\langle\rho^{\mathrm{SIDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}\rangle=0.3^{+0.9}_{-0.2}$ of its CDM counterparts, as well as shallower density profiles: $\langle\alpha_{\mathrm{SIDM}}\rangle=-1.4^{+0.4}_{-0.3}$ vs $\langle\alpha_{\mathrm{CDM}}\rangle=-1.7^{+0.3}_{-0.4}$. In WDM, the profiles change in the opposite direction than SIDM, being instead slightly denser, $\langle\rho^{\mathrm{WDM}}_{\mathrm{exsitu}}/\rho^{\mathrm{CDM}}_{\mathrm{exsitu}}\rangle=1.4^{+0.6}_{-0.6}$, and steeper, $\langle\alpha_{\mathrm{WDM}}\rangle=-2.0^{+0.1}_{-0.4}$ than the CDM counterparts. However, the changes in the WDM ex-situ density are minor compared to those observed in the SIDM model.
  • Beyond their density profiles, the velocity distribution of accreted stars also differs across models. This results in a larger number and prominence of overdensities in $r,v_{r}$ space (Fig. 3) and more tangentially supported orbits in SIDM than in CDM (Fig. 4). This amounts to a typical change in the velocity anisotropy of $\langle\Delta\beta\rangle=-0.2^{+0.1}_{-0.1}$ between SIDM and CDM in the inner $0.01R_{200}$ ($\approx 2\,\mathrm{kpc}$). The changing anisotropy could be used in conjunction with the properties of the GE/S in the Milky Way to place upper limits on the allowed SIDM cross-section. In other words, what is the maximum possible value of $\sigma_{\mathrm{SIDM}}$ that can accommodate highly radially supported ($\beta\approx 0.8$) remnant of GE/S-like events. This will be explored in upcoming work.
@@
Several assumptions and limitations are present in this study. First, we have considered simulations in which the predicted differences between dark matter models are exacerbated, i.e. no flat-density core formation through gas blowouts in dwarf galaxies. As shown in Forouhar Moreno et al. (2022), this process can mimic the effects of SIDM on the satellite population, and could therefore similarly affect the most massive stellar halo progenitors and thus its final properties. It nonetheless remains to be seen whether baryon-driven outflows are effective at forming flat density cores for massive GE/S-like galaxies at high redshifts. Second, our comparisons between dark matter models do not include observational uncertainties. This means that the differences we identified in this work may be difficult to observe in the real Universe. As such, mock observations should be made based on these simulations, to enable a fairer comparison to real data. Third, we have used relatively extreme variations in the nature of dark matter, and so follow-up work should consider versions that are more consistent with current observational constraints.
Regarding the last point, one may indeed wonder how our findings for the SIDM model would change if we had used a velocity-dependent cross-section. Based on Fig. 5, the most massive progenitor of the stellar haloes found in our sample typically reach $V_{\rm peak}\approx 100\,\mathrm{km}\,\mathrm{s}^{-1}$. We can therefore question whether dark matter haloes with $V_{\rm max}=100\,\mathrm{km}\,\mathrm{s}^{-1}$ have non-negligible velocity-dependent cross-sections, and hence affect their structure significantly. If so, differences in the most massive progenitors of the stellar halo will be similar to those discussed in this study, and hence result in significant differences in the kinematic and spatial distributions of ex-situ stars relative to CDM.
Several velocity-dependent cross-sections models have been used in the literature (e.g. Dooley et al., 2016; Correa, 2021; Correa et al., 2024). The models generally share similar trends at large and small relative velocities, which motivated by cluster- and dwarf-scale constraints, respectively require small and large cross-section values. However, there is up to an order of magnitude difference in the expected cross-section at $\Delta v\approx 100\,\mathrm{km}\,\mathrm{s}^{-1}$, owed in part to the relative lack of constraints at these scales. Hence, whether the most significant progenitor is expected to be affected when using a velocity-dependent cross-section depends on the assumed model. Taking as an example the model used in recent cosmological hydrodynamical simulations (e.g. Correa et al., 2024), and using the maximum 1D velocity dispersion of the halo as our reference relative velocity, the expected cross-section of haloes with $V_{\mathrm{max}}=100\,\mathrm{km}\,\mathrm{s}^{-1}$ is at least $\approx 5.25\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$.
A velocity-dependent cross-section would have other implications beyond the most massive progenitors. The high cross-section values for the lowest mass progenitor galaxies could trigger gravothermal collapse, making their central profiles denser than an NFW and hence more resilient to tides. Hence, the number and properties of tidal debris originating from them could significantly change relative to our SIDM model and CDM models. In terms of the Milky Way host dark matter halo, the low values of cross-section expected at its mass scale would likely mean that no flat density core forms. Compared to the constant cross-section we have used in this study, which does have a flat density core, the strength of the tidal field would be enhanced in the central regions.
@@
Despite these caveats, our findings highlight a new potential avenue to constrain the nature of dark matter, based on the dynamics and distribution of the accreted stellar halo of Milky Way-mass galaxies. One approach would be to measure the spatial distribution of the stellar halo around a large sample of MW-like galaxies, and compare it to the predicted distribution across different DM models. An alternative would be to leverage the phase-space information of GE/S, by identifying simulated MW-analogues that experienced similar events in their past and measuring the final velocity anisotropy of its remnant. As changing the value of the SIDM cross-section systematically affects the merging process, and hence the anisotropy $\beta$, comparing the predicted distributions of $\beta$ to the one observed for GE/S could hint at which SIDM cross-sections are compatible with the data.
@@
Beyond these two approaches, one can imagine using other observables which we have not considered and could be similarly affected, such as the two-point correlation function in phase-space, clustering in action space, and metallicity gradients. Given ongoing and upcoming efforts towards studying the stellar halo Milky Way and those of external Milky Way-mass haloes, this exciting prospect warrants further investigation.
@@ Acknowledgements
We thank the anonymous referee for their comments and suggestions, which have improved the presentation and clarify of this paper. VJFM acknowledges support by NWO through the Dark Universe Science Collaboration (OCENW.XL21.XL21.025) and the European Research Council (ERC) through the Advanced Investigator grant DMIDAS (GA 786910) and Consolidated Grant ST/T000244/1. ABL acknowledges support by the Italian Ministry for Universities (MUR) program “Dipartimenti di Eccellenza 2023-2027” within the Centro Bicocca di Cosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). AD is supported by a Royal Society University Research Fellowship and the Leverhulme Trust. AD and AF acknowledge support from the Science and Technology Facilities Council (STFC) [grant numbers ST/X001075/1, ST/T000244/1]. AF has been supported by a UKRI Future Leaders Fellowship (grant no MR/T042362/1) and a Wallenberg Acadamy Fellowship. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@ Appendix B Radial velocity vs radial distance of accreted stars
There is a clear example of a stellar halo, that of galaxy G4, which is noticeably asymmetrical about the $v_{r}=0$ line. The asymmetrical features are caused by the ongoing disruption of a massive satellite that was accreted at $t_{\mathrm{infall}}\approx 7.69~\mathrm{Gyr}$. The merger is deemed as complete for the SIDM simulation, as the galaxy is no longer found in the galaxy catalogues. For CDM and WDM, the companion is still found as a self-bound satellite galaxy, although only a small amount of central mass remains self-bound.
Omitting the example of G4 due to the ongoing disruption of a massive satellite, we observe the same trends as those discussed in the main text. The chevrons appear in greater numbers in SIDM, and are located further out than those in CDM and WDM. However, the extend of the differences between dark matter models is highly dependent on how long ago the last major merger happened. Galaxies which had more recent major mergers have chevrons that enhance the differences found across dark matter models. The presence of chevrons is simply due to the fact that not enough time has passed for them to phase-mix into the background. In contrast, the galaxies that experienced their last major merger earlier on have largely featureless distributions of $v_{r}$ vs $r$, blurring the differences in initial chevron populations due to different dark matter models. The trend we observe between the ease of identifying differences across dark matter models based on chevrons with how long ago the last major merger happened highlights that the most constraining galaxies are likely to be those that form late.
Figure 8: Radial velocity of accreted stellar particles, as a function of their radial distance to the centre of their host galaxies. The colour of each particle encodes the local ex-situ stellar density. This is shown for the remainder of the haloes not shown in Fig. 3, with their their matched CDM, WDM and SIDM versions placed from left to right. The vertical ordering reflects the $M_{200}(z=0)$ ranking, with decreasing mass towards the lower panels. The infall time of the satellite galaxy that was involved in the last major merger for each of the galaxies in this sample is indicated on the right-hand side of the corresponding row.